# 2-Cats: 2D Copula Approximating Transforms

### v3

@@ Abstract
Copulas are powerful statistical tools for capturing dependencies across multiple data dimensions. Applying Copulas involves estimating independent marginals, a straightforward task, followed by the much more challenging task of determining a single copulating function, $C$, that links these marginals. For bivariate data, a copula takes the form of a two-increasing function $C:(u,v)\in\mathbb{I}^{2}\rightarrow\mathbb{I}$, where $\mathbb{I}=[0,1]$. In this paper, we propose 2-Cats, a Neural Network (NN) model that learns two-dimensional Copulas while preserving their key properties, without relying on specific Copula families (e.g., Archimedean). Furthermore, we introduce a training strategy inspired by the literature on Physics-Informed Neural Networks and Sobolev Training. Our proposed method exhibits superior performance compared to the state-of-the-art across various datasets while maintaining the fundamental mathematical properties of a Copula.
@@ 1 Introduction
Modeling univariate data is relatively straightforward for several reasons. Firstly, there exists a wide range of probability distributions suitable for different types of data, including Gaussian, Beta, Log-Normal, Gamma, and Poisson. This set is expanded by mixture models that use the basic distributions as components and can model multimodality and other behavior. Secondly, visual tools such as histograms or Kernel Density Estimators (KDE) provide valuable assistance in selecting the appropriate distribution. Lastly, univariate models typically involve few parameters and may even have exact solutions, simplifying the modeling process.
@@
However, the process becomes more challenging when modeling multivariate data and obtaining a joint probability distribution. Firstly, only a few classes of multivariate probability distributions are available, with the primary exceptions being the multivariate Gaussian, Elliptical, and Dirichlet distributions. Secondly, simultaneously identifying the dependencies among multiple variables is not easy. Distinguishing direct dependencies from conditional independence based solely on empirical data is a highly complex task.
@@
In 1959, Abe Sklar formalized the idea of Copulas [Sklar, 1959, 1996]. This statistical tool allows us to model a multivariate random variable of dimension $d$ by learning its cumulative multivariate distribution function (CDF) using $d$ independent marginal CDFs and a single additional Copulating function $C$. A bivariate example may separately identify one marginal as a Log-Normal distribution and the other as a Beta distribution. Subsequently, the Copula model links these two marginal distributions.
@@
However, the Copula approach to modeling multivariate data faced a significant limitation until recently. Our model choices were confined to only a handful of closed forms for the $C$ Copulas, such as the Gaussian, Frank, or Clayton Copulas [Czado and Nagler, 2022, Größer and Okhrin, 2022]. Unfortunately, this approach using closed-form Copula models proved inadequate for accurately capturing the complex dependencies in real data. The limited parameterization of these Copulas prevented them from fully representing the intricate relationships between variables.
@@
This situation has changed recently. Neural Networks (NNs) are well-known for their universal ability to approximate any function. Consequently, researchers have started exploring NNs as alternatives to closed forms for $C$ [Chilinski and Silva, 2020, Hirt et al., 2019, Tagasovska et al., 2023, Janke et al., 2021, Ling et al., 2020, Ng et al., 2021]. However, a limitation of several NN models (not only for Copulas) is that they neglect the importance of maintaining the fundamental mathematical properties of their domain of study.
@@
For Copulas, only three properties define $C$ [Nelsen, 2006]: P1: $C:(u,v)\in\mathbb{I}^{2}\mapsto\mathbb{I}$, where $\mathbb{I}=[0,1]$.
P2: $C$ is grounded. That is, $C(0,v)=0$, $C(v,0)=0$. Moreover, $C(1,v)=v$ and $C(u,1)=u$.
P3: For any $0\leq u_{1}<u_{2}\leq 1$ and $0\leq v_{1}<v_{2}\leq 1$, we have that the volume of $C$ is: $V_{C}(u_{1},u_{2},v_{1},v_{2})\equiv C(u_{2},v_{2})-C(u_{2},v_{1})-C(u_{1},v_{2})+C(u_{1},y_{1})\geq 0$.
Our model, 2-Cats, is a NN approach that meets such properties. The formal proofs are in Section 3.2
[…]
@@
In this paper, we propose 2-Cats (2D Copula Approximating Transforms). To understand our approach, let $H_{\theta}(u,v)$ represent our copula model (or Hypothesis). Let $G:\mathbb{R}^{2}\mapsto\mathbb{I}$ be any bivariate CDF with $\mathbb{R}^{2}$ as support. Examples are the bivariate Normal or bivariate Logistic. We connect $H_{\bm{\theta}}$ with $G$ as follows:
@@
Here, $m_{\bm{\theta}}(u,v)$ is any NN that outputs a positive number and $z(x)$ is the Logit function that is both monotonic and maps $[0,1]$ to [,][$\infty$,$\infty$][ - ∞ , ∞ ], i.e., $z:\mathbb{I}\mapsto\mathbb{R}$. Here, the function $t_{v}(u)$ is monotonically increasing concerning $u$ because of the positive NN output; similarly, $t_{u}(v)$ is monotonic on $v$.
The monotonic properties of all transformations ensure that our output $H_{\bm{\theta}}(u,v)$ is monotonic for both $u$ and $v$. Let $z_{u}=z(t_{v}(u))$ and $z_{v}=z(t_{u}(v))$ to simplify notation and $(U,V)$ a random vector with distribution function $H_{\bm{\theta}}(u,v)$. Then, the probability transform below is valid [Casella and Berger, 2001, section 4.3]:
$H_{\bm{\theta}}(u,v)$ $=\Pr\big{[}U\leq u,V\leq v\big{]}$
$=\Pr\big{[}t^{-1}_{v}(U)\leq t^{-1}_{v}(u),t^{-1}_{v}(V)\leq t^{-1}_{v}(v)\big{]};$
$=\Pr\big{[}Z_{u}\leq z_{u},Z_{v}\leq z_{v}\big{]}$
As we prove in Section 3.2, this guarantees that $H_{\bm{\theta}}(u,v)$ is a Copula and meets the P1-P3 properties.
There have been several recent NN Copula/CDF-based models [Chilinski and Silva, 2020, Ling et al., 2020, Ng et al., 2021] and non-NN models Nagler et al. [2017], Bakam and Pommeret [2023], Mukhopadhyay and Parzen [2020]. We compare our 2-Cats model with these alternatives and our results in Section 5 show that the 2-Cats performance (in negative log-likelihood) is better or statistically tied to the results from the state-of-the-art.
A second contribution of our work is that 2-Cats models are trained similarly to Physics Informed Neural Networks (PINNs) [Karniadakis et al., 2021, Lu et al., 2021]. We employ Sobolev training [Czarnecki et al., 2017] to minimize losses concerning an objective function and its derivatives. Ours is the first NN-based Copula work to do so and it bridges two research areas.
In summary, our contributions can be outlined as follows:
  • We introduce the 2-Cats model, a novel neural network-based Copula that not only satisfies all the essential requirements of a Copula (refer to Section 3.2), but also demonstrates empirical accuracy as an effective approximator for various 2-D Copulas.
  • Beyond meeting the fundamental properties of a Copula, our model enables the estimation of the Jacobian and Hessian of the Copula through symbolic differentiation. This capability allows us to assess several crucial aspects: the likelihood of the Copula (or the pseudo-likelihood of the data, defined as the second derivative); the conditional empirical distributions of the Copula (first derivative); the inverse of the previous function, inverse of the conditional, useful for sampling; and, the cumulative function (the Copula itself).
  • We are the first to apply Sobolev training [Czarnecki et al., 2017] in NN-based Copula methodologies. Our approach involves the introduction of a simple empirical estimator for the first derivative of a Copula, seamlessly integrating it into our training framework.
@@ 2 Related Work
After being introduced by Sklar [Sklar, 1959, 1996], Copulas have emerged as valuable tools across various domains [Größer and Okhrin, 2022, Nelsen, 2006], including engineering [Salvadori and De Michele, 2004], geosciences [Modiri et al., 2020, Yang et al., 2019], and finance [Cherubini et al., 2004, Salvadori and De Michele, 2004]. Recently, Copulas have gained attention from the machine learning (ML) community, particularly with the advent of Deep Learning methods for Copula estimation and their utilization in generative models [Chilinski and Silva, 2020, Hirt et al., 2019, Tagasovska et al., 2023, Ling et al., 2020, Janke et al., 2021, Ng et al., 2021].
@@
Our proposed approach aligns closely with the emerging literature in this domain but makes significant advancements on two fronts. This paper primarily focuses on developing an NN-based Copula model that adheres to the fundamental properties of a Copula. Two methods sharing similar objectives are discussed in [Ling et al., 2020, Ng et al., 2021] and [Chilinski and Silva, 2020]. However, our approach diverges from these prior works by not confining our method to Archimedean Copulas, as seen in [Ling et al., 2020] and [Ng et al., 2021].
@@
As Chilinski and Silva [Chilinski and Silva, 2020], we begin by estimating densities through CDF estimations followed by automatic differentiation to acquire densities. However, our primary differentiation lies in our emphasis on fulfilling the requisite properties of a copula, an aspect not prioritized by the authors of that study.
@@
To assess the effectiveness of our approach, we conduct comparative analyses against previous proposals, including the method presented by Janke et al. [Janke et al., 2021], proposed as a generative model. Furthermore, 2-Cats does not presume a fixed density function for copula estimation; instead, it is guided by the neural network (NN). Hence, our work shares similarities with non-parametric copula estimators used as baselinesNagler et al. [2017], Mukhopadhyay and Parzen [2020], Bakam and Pommeret [2023]. Ultimately, all these methods function as baseline approaches against which we evaluate and demonstrate the superiority of our proposed approach.
@@
For training our models, we utilize Sobolev training as a regularizer. Sobolev training is tailored to estimate neural networks (NNs) by incorporating losses that not only address errors toward the primary objective but also consider errors toward their derivatives. This approach ensures a more comprehensive and accurate training of the NNs. A similar concept is explored in Physics Informed Neural Networks (PINNs) [Karniadakis et al., 2021, Kobyzev et al., 2020], where NNs incorporate derivative information from physical models. This derivative-aware training enables PINNs to better capture complex physics-related relationships.
@@
In summary, our objective is to develop an NN-based Copula model that respects the mathematical properties of Copulas and benefits from considering derivative information. Consequently, although our focus is on 2D Copulas, the Pair Copula Decomposition (PCC) [Aas et al., 2009, Czado, 2010] that requires derivative information and is used to model $d$-dimensional data is valid for our models [Czado and Nagler, 2022]. Due to space constraints, we leave the evaluation of such a decomposition for future work.
@@
Our work also aligns with the literature on monotonic networks [Sill, 1997, Daniels and Velikova, 2010, Wehenkel and Louppe, 2019]. The concept of using integral transforms to build monotonic models was introduced by Wehenkel and Louppe [2019]. This idea is currently evolving in the form of Integral Networks [Kortvelesy, 2023, Liu, 2023].
3 Properties and Proofs
We now revisit the discussion on the 2-Cats model as introduced in Section 1. Specifically, we will elaborate on the model’s implementation and rigorously establish the validity of 2-Cats as a copula.
3.1 Implementation Details
@@
One of the building blocks of 2-Cats is the NN $m_{\bm{\theta}}(u,v)$. In particular, we require that this NN outputs only positive numbers and that its integral captures a monotonic function. We achieve this by employing Exponential Linear Units (ELU) plus one activation (ELU+1) in every layer as is the Unconstrained Monotonic Network defined by  Wehenkel and Louppe [2019].
@@
Let us consider the computation details of $t_{v}(u)$, with the understanding that the process for $t_{u}(v)$ is analogous. In our implementation, we adopt Cumulative Trapezoid integration. Initially, we divide the range $[0,1]$ into $200$ equally spaced intervals. When computing $t_{v}(u)$, the value of $u$ is inserted into this interval while preserving the order. Subsequently, the $m_{\bm{\theta}}(x,v)$ NN is evaluated for the $201$ corresponding values of parameter $x$. Trapezoidal integration is then applied to calculate the integral value at $u$ (for the numerator) and $1$ (for the denominator). Finally, when $u=0$, it follows that $t_{v}(0)=0$.
@@
To compute the Jacobian and Hessian of $H_{\bm{\theta}}$, as is done in PINNs [Karniadakis et al., 2021, Lu et al., 2021], we rely on symbolic computation from modern differentiable computing frameworks such as Jax and Pytorch for this case. In other words, $\frac{\partial H_{\bm{\theta}}(u,v)}{\partial u}$ and $\frac{\partial H_{\bm{\theta}}(u,v)}{\partial v}$ are estimated via symbolic differentiation via the Jacobian [^1] https://jax.readthedocs.io/en/latest/_autosummary/jax.jacobian.html, while $\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{\partial u\partial v}$ is estimated via symbolic differentiation using the Hessian [^2] https://jax.readthedocs.io/en/latest/_autosummary/jax.hessian.html. Our Cumulative Trapezoidal approach for integrating does not affect this symbolic differentiation as the intervals are fixed, and the Trapezoidal rule is comprised of basic arithmetic operations (sums, subtractions, multiplications, and division).
3.2 Proofs
We present a detailed proof that 2-Cats is a Copula, thereby inheriting all Copula properties as outlined in [Nelsen, 2006, Chapter 2]. For details on how to sample from 2-Cats, please refer to Appendix A.
@@ 3.2 Proofs
Remind that $z_{u}=z\big{(}t_{v}(u)\big{)}$ and $z_{v}=z\big{(}t_{u}(v)\big{)}$. By construction, we have $(t_{v}(u),t_{u}(v))\in\mathbb{I}^{2}$. These transforms are also able to cover the entire range of $\mathbb{I}^{2}$ (when $u=0$ we have $t_{v}(0)=0$, when $u=1$ we have $t_{v}(1)=1$, the same goes for the $t_{u}(v)$ transform). The function $z$ maps the domain $\mathbb{I}^{2}$ into $\mathbb{R}^{2}$. Given that the bivariate CDFs work in this domain, we have that: $H_{\bm{\theta}}:(u,v)\in\mathbb{I}^{2}\mapsto\mathbb{I}$, where $\mathbb{I}=[0,1]$. ∎
Theorem P2.
$H_{\bm{\theta}}$ is grounded. That is, $H_{\bm{\theta}}(0,v)=0$, $C(v,0)=0$. Moreover, $H_{\bm{\theta}}(1,v)=v$ and $H_{\bm{\theta}}(u,1)=u$.
Proof.
Consider the value of $H_{\bm{\theta}}(u,0)$ when $v=0$. As $t_{u}(0)=0$, we have $z(t_{u}(0))=z(0)=\lim_{x\rightarrow 0}\text{Logit}(x)=-\infty$ and hence $H_{\bm{\theta}}(u,0)=G(z_{u},-\infty)$. Given that $G$ is a bivariate CDF, $G(z_{u},-\infty)=\lim_{w\rightarrow-\infty}G(z_{u},w)=0$ for any $z_{u}$.
We show now that $H_{\bm{\theta}}(u,1)=u$. As $t_{u}(1)=1$, we have
$H_{\bm{\theta}}(u,1)$ $=G(z(t_{1}(u)),z(t_{u}(1)))=G(z(t_{1}(u)),z(1))$
$=G(z_{u},\infty)=\Pr[Z_{u}\leq z_{u}]=\Pr[t_{v}(U)\leq z_{u}]$
$=\Pr[U\leq{t_{v}^{-1}}(z_{u})]=\Pr[U\leq u]=u$
where $Z_{u}$ is the random variable associated with $z_{u}$. The above relationships are also valid because $t_{v}$ is monotonic and, consequently, invertible. $\Pr[U\leq u]=u$ is by definition as $U\leavevmode\nobreak\ U(0,1)$. Analogously, we prove that $H_{\bm{\theta}}(0,v)=0$ and $H_{\bm{\theta}}(1,v)=v$, concluding the proof. ∎
@@
This is a straightforward consequence that $G$ is a bivariate cumulative distribution function and the monotonicity of the $z$, $t_{u}$ and $t_{v}$ transforms. The same transform we explored above for the single variate case is valid for multiple variables. That is, fact that $z_{u}=z(t_{v}(u))$ and $z_{v}=z(t_{u}(v))$ are monotonic, one-to-one, guarantees that $G(z_{u},z_{v})$ defines a valid probability transform to a new CDF [Casella and Berger, 2001, section 4.3].
$\Pr\big{[}U\leq u,V\leq v\big{]}=\Pr\big{[}Z_{u}\leq z_{u},Z_{v}\leq z_{v}\big{]}.$

### v4

@@ Abstract
Copulas are powerful statistical tools for capturing dependencies across data dimensions. Applying Copulas involves estimating independent marginals, a straightforward task, followed by the much more challenging task of determining a single copulating function, $C$, that links these marginals. For bivariate data, a copula takes the form of a two-increasing function $C:(u,v)\in\mathbb{I}^{2}\rightarrow\mathbb{I}$, where $\mathbb{I}=[0,1]$. This paper proposes 2-Cats, a Neural Network (NN) model that learns two-dimensional Copulas without relying on specific Copula families (e.g., Archimedean). Furthermore, via both theoretical properties of the model and a Lagrangian training approach, we show that 2-Cats meets the desiderata of Copula properties. Moreover, inspired by the literature on Physics-Informed Neural Networks and Sobolev Training, we further extend our training strategy to learn not only the output of a Copula but also its derivatives. Our proposed method exhibits superior performance compared to the state-of-the-art across various datasets while respecting (provably for most and approximately for a single other) properties of C.
@@ 1 Introduction
Modeling univariate data is relatively straightforward for several reasons. Firstly, a wide range of probability distributions exist that are suitable for different data types, including Gaussian, Beta, Log-Normal, Gamma, and Poisson. This set is expanded by mixture models that use the elemental distributions as components and can model multimodality and other behavior. Secondly, visual tools such as histograms or Kernel Density Estimators (KDE) provide valuable assistance in selecting the appropriate distribution. Lastly, univariate models typically involve few parameters and may have exact solutions, simplifying the modeling process.
@@
However, the process becomes more challenging when modeling multivariate data and obtaining a joint probability distribution. Firstly, only a few classes of multivariate probability distributions are available, with the primary exceptions being the multivariate Gaussian, Elliptical, and Dirichlet distributions. Secondly, simultaneously identifying the dependencies via conditional distributions based solely on empirical data is highly complex.
@@
In 1959, Abe Sklar formalized the idea of Copulas [61, 60]. This statistical tool allows us to model a multivariate random variable of dimension $d$ by learning its cumulative multivariate distribution function (CDF) using $d$ independent marginal CDFs and a single additional Copulating function $C$. A bivariate example may separately identify one marginal as a Log-Normal distribution and the other as a Beta distribution. Subsequently, the Copula model links these two marginal distributions.
@@
However, the Copula approach to modeling multivariate data faced a significant limitation until recently. Our model choices were confined to only a handful of closed forms for the $C$ Copulas, such as the Gaussian, Frank, or Clayton Copulas [16, 25]. Unfortunately, this approach using closed-form Copula models proved inadequate for accurately capturing the complex dependencies in real data. The limited parameterization of these Copulas prevented them from fully representing the intricate relationships between variables.
@@
This situation has changed recently. Neural Networks (NNs) are well-known for their universal ability to approximate any function. Consequently, researchers have started exploring NNs as alternatives to closed forms for $C$ [13, 27, 66, 31, 38, 51]. However, a limitation of NNs is that they neglect the importance of maintaining the fundamental mathematical properties of their domain of study.
@@
For Copulas, only three properties define $C$ [50]: P1: $C:(u,v)\in\mathbb{I}^{2}\mapsto\mathbb{I}$, where $\mathbb{I}=[0,1]$; P2: For any $0\leq u_{1}<u_{2}\leq 1$ and $0\leq v_{1}<v_{2}\leq 1$, we have that the volume of $C$ is: $V_{C}(u_{1},u_{2},v_{1},v_{2})\equiv C(u_{2},v_{2})-C(u_{2},v_{1})-C(u_{1},v_{2})+C(u_{1},y_{1})\geq 0$; and, P3: $C$ is grounded. That is, $C(0,v)=0$, $C(v,0)=0$. Moreover, $C(1,v)=v$ and $C(u,1)=u$.
This paper proposes 2-Cats (2D Copula Approximating Transforms) as a NN approach that meets the above desiderata. How the model meets P1 to P3 is discussed in Section 3. P1 and P2 are natural consequences of the model, whereas P3 is guaranteed via Lagragian Optimization.
@@
To understand our approach, let $H_{\theta}(u,v)$ represent our model (or Hypothesis). Let $G:\mathbb{R}^{2}\mapsto\mathbb{I}$ be any bivariate CDF with $\mathbb{R}^{2}$ as support. We connect $H_{\bm{\theta}}$ with $G$ as follows:
@@
Here, $m_{\bm{\theta}}(u,v)$ is any NN that outputs a positive number and $z(x)$ is the Logit function that is both monotonic and maps $[0,1]$ to [,][$\infty$,$\infty$][ - ∞ , ∞ ], i.e., $z:\mathbb{I}\mapsto\mathbb{R}$. Here, the function $t_{v}(u)$ monotonically increases concerning $u$ because of the positive NN output; similarly, $t_{u}(v)$ is monotonic on $v$.
The monotonic properties of all transformations ensure that our output $H_{\bm{\theta}}(u,v)$ is monotonic for both $u$ and $v$. Let $z_{u}=z(t_{v}(u))$ and $z_{v}=z(t_{u}(v))$ to simplify notation and $(U,V)$ a random vector with distribution function $H_{\bm{\theta}}(u,v)$. Then, the probability transform below is valid [11, section 4.3]:
$H_{\bm{\theta}}(u,v)$ $=\Pr\big{[}Z_{u}\leq z_{u},Z_{v}\leq z_{v}\big{]}=\Pr\big{[}t^{-1}_{v}(U)\leq t^{-1}_{v}(u),t^{-1}_{v}(V)\leq t^{-1}_{v}(v)\big{]}$
$=\Pr\big{[}U\leq u,V\leq v\big{]}$ (2)
To ensure that our model derivatives also approximate the derivatives of $C$, a second contribution of our work is that 2-Cats models are trained similarly to Physics Informed Neural Networks (PINNs) [32, 42] by employing Sobolev training [17] approximate Copula derivatives.
There have been several recent NN [13, 38, 51] and non-NN models [49, 5, 47] Copula-like models. We compare our 2-Cats model with these alternatives, and our results show that the 2-Cats performance (in negative log-likelihood) is better or statistically tied to baselines. Our contributions are:
  • We introduce the 2-Cats model, a novel NN Copula approximation. Unlike other NN approaches, we focus on satisfying (either provably or via constraints) the essential requirements of a Copula. We also demonstrate the empirical accuracy of 2-Cats;
  • We are the first to apply Sobolev training [17] in NN-based Copula methodologies. Our approach involves the introduction of a simple empirical estimator for the first derivative of a Copula, which is seamlessly integrated into our training framework.
@@ 2 Related Work
After being introduced by Sklar [61, 60], Copulas have emerged as valuable tools across various domains [25, 50], including engineering [56], geosciences [46, 72], and finance [12, 56]. Recently, Copulas have gained attention from the ML community, particularly with the advent of DL methods for Copula estimation and their utilization in generative models [13, 27, 66, 38, 31, 51].
@@
Our proposed approach aligns closely with the emerging literature in this domain but makes significant advancements on two fronts. This paper primarily focuses on developing an NN-based Copula model that adheres to the fundamental properties of a Copula. Two methods sharing similar objectives are discussed in [38, 51] and [13]. However, our approach diverges from these prior works by not confining our method to Archimedean Copulas, as seen in [38] and [51].
@@
As Chilinski and Silva [13], we begin by estimating densities through CDF estimations followed by automatic differentiation to acquire densities. However, our primary difference lies in our emphasis on relating 2-Cats to a Copula, an aspect not prioritized by the authors of that study.
@@
To assess the effectiveness of our approach, we conduct comparative analyses against previous proposals, including the method presented by Janke et al. [31], proposed as a generative model. Furthermore, 2-Cats does not presume a fixed density function for copula estimation; instead, it is guided by the neural network (NN). Hence, our work shares similarities with non-parametric copula estimators used as baselines[49, 47, 5]. Ultimately, all these methods function as baseline approaches against which we evaluate and demonstrate the superiority of our proposed approach.
@@
For training our models, we utilize Sobolev training as a regularizer. Sobolev training is tailored to estimate neural networks (NNs) by incorporating losses that not only address errors toward the primary objective but also consider errors toward their derivatives. This approach ensures a more comprehensive and accurate training of the NNs. A similar concept is explored in Physics Informed Neural Networks (PINNs) [32, 35], where NNs incorporate derivative information from physical models. This approach enables PINNs to model complex physics-related relationships.
@@
In summary, our objective is to develop an NN-based Copula model that respects the mathematical properties of Copulas and benefits from considering derivative information. Consequently, although our focus is on 2D Copulas, the Pair Copula Decomposition (PCC) [1, 15] that requires derivative information and is used to model $d$-dimensional data is valid for our models [16]. Due to space constraints, we leave the evaluation of such a decomposition for future work.
@@
Our work also aligns with the literature on monotonic networks [58, 18, 71] and Integral Networks [36, 39]. The concept of using integral transforms to build monotonic models was introduced by [71].
3 On the Desiderata of a Copula and 2-Cats Properties
We now revisit the discussion on the 2-Cats model as introduced in Section 1.
@@ 3 On the Desiderata of a Copula and 2-Cats Properties
One of the building blocks of 2-Cats is the NN $m_{\bm{\theta}}(u,v)$. In particular, we require that this NN outputs only positive numbers so that its integral captures a monotonic function. We achieve this by employing Exponential Linear Units (ELU) plus one activation (ELU+1) in every layer.
@@
Let us consider the computation details of $t_{v}(u)$, with the understanding that the process for $t_{u}(v)$ is analogous. In our implementation, we adopt Cumulative Trapezoid integration. Initially, we divide the range $[0,1]$ into $200$ equally spaced intervals. When computing $t_{v}(u)$, the value of $u$ is inserted into this interval while preserving the order. Subsequently, the $m_{\bm{\theta}}(x,v)$ NN is evaluated for the $201$ corresponding values of parameter $x$. Trapezoidal integration is then applied to calculate the integral value at $u$ (for the numerator) and $1$ (for the denominator). Finally, when $u=0$, $t_{v}(0)=0$.
@@
To compute the Jacobian and Hessian of $H_{\bm{\theta}}$, as is done in PINNs [32, 42], we rely on symbolic computation from modern differentiable computing frameworks such as Jax and Pytorch for this case. In other words, $\frac{\partial H_{\bm{\theta}}(u,v)}{\partial u}$ and $\frac{\partial H_{\bm{\theta}}(u,v)}{\partial v}$ are estimated via symbolic differentiation via the Jacobian [^1] https://jax.readthedocs.io/en/latest/_autosummary/jax.jacobian.html, while $\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{\partial u\partial v}$ is estimated via symbolic differentiation using the Hessian [^2] https://jax.readthedocs.io/en/latest/_autosummary/jax.hessian.html.
Now, let us present our main properties. Appendix B and  D complements these properties.
@@
Remind that $z_{u}=z\big{(}t_{v}(u)\big{)}$ and $z_{v}=z\big{(}t_{u}(v)\big{)}$. Also, $(t_{v}(u),t_{u}(v))\in\mathbb{I}^{2}$. These transforms are also able to cover the entire range of $\mathbb{I}^{2}$ (when $u=0$ we have $t_{v}(0)=0$, when $u=1$ we have $t_{v}(1)=1$, the same goes for the $t_{u}(v)$ transform). The function $z$ maps the domain $\mathbb{I}^{2}$ into $\mathbb{R}^{2}$. Given that the bivariate CDFs work in this domain: $H_{\bm{\theta}}:(u,v)\in\mathbb{I}^{2}\mapsto\mathbb{I}$, where $\mathbb{I}=[0,1]$. ∎
@@
This is a straightforward consequence of $G$ being a bivariate cumulative distribution function and the monotonicity of the $z$, $t_{u}$, and $t_{v}$ transforms. The same transform we explored above for the single variate case is valid for multiple variables. That is, fact that $z_{u}=z(t_{v}(u))$ and $z_{v}=z(t_{u}(v))$ are monotonic guarantees that $G(z_{u},z_{v})$ defines a valid transform Eqn (2).